OpenStudy (anonymous):

Find the intersection point of the two curves below: y = 4sinx y = e^{4x}

OpenStudy (anonymous):

Can this be done algebraically?

OpenStudy (anonymous):

solve \(4 \sin x=e^{4x}\)

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

How can you get the x-term by itself?

OpenStudy (anonymous):

so is that 2 graphs y=sinx AND y = e^4x?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

4sinx = e^(4x)

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

o wth that looks confusing

OpenStudy (anonymous):

How you tried graphing the two functions to see where they cross?

OpenStudy (anonymous):

Ok lets say I am not allowed to use a graphing calculator.

OpenStudy (anonymous):

ok lets take ln on both sides to get ln(4sinx) = 4x on the left side we can right ln 4 + ln(sinx) = 4x

OpenStudy (anonymous):

The you can use pencil and paper which is much better anyway.

OpenStudy (turingtest):

I really doubt there is an analytic way to do this best using some approximation technique like newton-ralphson or somethin

OpenStudy (anonymous):

@markriggs Yes, but that requires me to get the intersection points. I have to find the area of the region b/ween y = 4sinx and y =e^{4x} bounded by x = 0 and x = pi/2. That is the full question

OpenStudy (anonymous):

And for me to do that, I need to find the intersection points in the MIDST of these boundaries

OpenStudy (anonymous):

The two functions do not cross on the interval (0,pi/2).

OpenStudy (anonymous):

Really?

OpenStudy (turingtest):

really

OpenStudy (anonymous):

No I am pretty sure they do

OpenStudy (turingtest):

sinx starts at 0 e^(4x) starts at 1 and grows faster they do not intersect in that interval

OpenStudy (anonymous):

Wow

OpenStudy (anonymous):

markriggs, that was great. I was completely doubting that because I never really ended up with a situation like that

OpenStudy (anonymous):

Ok I think that is about it. I can work out the area. Next time I will check to make sure with graphing software if they ever intersect

OpenStudy (anonymous):

Sin(0)=0 e^(4*0)=1 Sin(pi/2)=1 e^2Pi=535 You really should draw a picture first..

OpenStudy (anonymous):

No I did. But I drew it incorrectly.

OpenStudy (anonymous):

Ok thanks all